# ViSymRe: Vision Multimodal Symbolic Regression

### v3

@@ Abstract
Extracting interpretable equations from observational datasets to describe complex natural phenomena is one of the core goals of artificial intelligence. This field is known as symbolic regression (SR). In recent years, Transformer-based paradigms have become a new trend in SR, addressing the well-known problem of inefficient inference. However, the modal heterogeneity between datasets and equations often hinders the convergence and generalization of these models. In this paper, we propose ViSymRe, a Vision Symbolic Regression framework that explores the positive role of visual modality in enhancing the performance of Transformer-based SR paradigms. To extend ViSymRe to high-dimensional scenarios, we present Multi-View Random Slicing (MVRS). By projecting multivariate equations into 2-D space using random affine transformations, MVRS avoids common defects in high-dimensional visualization, such as variable degradation, non-linear interaction missing, and exponentially increasing sampling complexity. To ensure dataset-only inference, we design a dual-visual pipeline that reconstructs visual features directly from the datasets via an auxiliary Visual Decoder and suppresses reconstruction noise via biased attention weights. Ablation studies demonstrate the positive contribution of visual modality to improving model convergence and enhancing various SR metrics. Furthermore, evaluation results on mainstream benchmarks indicate that ViSymRe achieves competitive performance compared to baselines, particularly in low-complexity and rapid-inference scenarios.
@@ 1 Introduction
Recent advances in SR have introduced Transformer architectures [3, 62, 14, 64, 5], reformulating the SR problem as a dataset-to-equation translation task, similar to machine translation models. Different from traditional GP and DL-based models, Transformer-based paradigms leverage the attention mechanism to capture dependencies within datasets, enabling direct equation generation without iteration. While improving deployment efficiency, these models often fail to adequately account for the modal gap between datasets and equations, which increases convergence difficulty. MMSR [30] pioneered multimodal learning in SR. It considers the fusion representation of the datasets and equations as direct inputs during training, and introduces multimodal techniques such as Cross-Attention feature fusion and alignment loss, demonstrating excellent performance compared to dataset-only baselines across several scenarios. Our solution further advances multimodal learning by incorporating visual modality, a readily available and crucial resource, but has been overlooked in previous research.
Currently, visual multimodal models have proven effective for enhancing various machine learning tasks, especially in machine translation. [11, 54]. The rationale is intuitive. For example, in machine translation, the English word “Dog” and the German word “Hund” share the same underlying visual perception, despite originating from different languages. Analogously, in SR, although datasets and equations present different forms, they correspond to the same geometric topology.
@@
Based on MVRS, we propose ViSymRe, a Transformer-based Vision Symbolic Regression framework. While MVRS enables low-cost visualization, its application is limited by the analytical equations, a condition that holds only during training. To achieve dataset-only inference, ViSymRe employs an auxiliary Visual Decoder to predict visual features directly from dataset embeddings, thereby eliminating the need for real visual input. Diverging from traditional pixel-level reconstruction or direct feature regression, we incorporate a Vector Quantization mechanism to project visual features into a discrete and finite space [63]. Specifically, ViSymRe maintains a learnable Codebook that discretizes the continuous visual feature space into a finite set of codewords. The Visual Decoder performs multi-head classification conditioned on the dataset to predict the indices of codewords corresponding to the visual features. By transforming the visual reconstruction task from regression to classification, we avoid the common mean collapse issue in high-dimensional regression [36]. In essence, ViSymRe establishes a dual-visual pipeline: a real pipeline and a virtual pipeline. The former supervises the convergence of the Visual Encoder to ensure the validity of visual features, while the latter enables dataset-only inference.
@@
    We propose ViSymRe, a Transformer-based framework that integrates the visual modality into SR and extends its applicability to high dimensions. Extensive evaluations demonstrate that the proposed architecture is robust across various scenarios, particularly showing strong potential for low-complexity and rapid-inference tasks.
@@
    We propose Multi-View Random Slicing (MVRS), a visualization strategy to support large-scale pre-training. MVRS projects multivariate equations into 2-D space without causing variable degeneracy or exponential sampling complexity.
@@
    We construct a dual-visual pipeline that introduces virtual vision to reconcile conflicts between multimodal training and dataset-only inference. By employing a Biased Cross-Attention feature fusion feature module to suppress attention weights assigned to noisy components in the virtual vision, the two vision pipelines are ensured to converge synchronously.
@@ 2 Motivation
Transformer-based paradigms have become a new trend for SR. They enable one-time inference, thus avoiding expensive iterative search. However, adapting the machine translation paradigms to SR remains challenging. A key difficulty lies in the modal heterogeneity between the datasets and equations, which limits the models’ convergence. Multimodal learning applied to SR has been preliminarily explored in [30], ensuring modal homogeneity by treating the equations as input. Our solution is to introduce the visual modality as a bridge to alleviate this limitation. The dataset-only paradigm requires models to generate equations from local, fragmented data points, whereas the multimodal paradigm provides a global perspective.
@@ 3 Related work
Overall, Transformer-based models offer an advantage in inference efficiency, avoiding the iterative search of GP and DL-based models by leveraging priors learned during training. However, these models face challenges in out-of-distribution generalization. To bridge this gap, some rectification strategies, such as TPSR [55], SNR [32], and SR-GPT [29], have been proposed to offer feedback about fitting losses and complexity to generation policies. These correction strategies are crucial for enhancing the generalization of Transformer-based models, significantly broadening their application scenarios at low cost.
@@ 4.2 Multi-view random slicing
Figure 2: Visualization results comparison of four visualization methods for the interaction term $x_{1}(x_{2}-x_{3})$. The proposed method (a) preserves the saddle shape via random isotropic slicing. Marginal projection (b) suffers from feature collapse, degenerating the non-linear interaction into a misleading linear response. Traditional methods (c,d): PCA exhibits information loss due to projection overlap, while t-SNE fails to reconstruct the manifold continuity, resulting in topological tearing even with dense sampling. Notably, MVRS requires only $N\cdot d$ points, while PCA and t-SNE require $N^{d}$ points. The computational burden is a key reason why traditional methods cannot support large-scale pre-training.
@@ 4.4 Visual Decoder
The Visual Decoder establishes a mapping between the dataset and Codebook, enabling dataset-only inference. This module operates essentially as a parallel classifier. We initialize $M$ learnable classification tokens $[CLS]s$, corresponding to the $M$ visual features output by the Visual Encoder. $[CLS]s$ aggregates with the dataset features via a Cross-Attention module to generate $M$ classification vectors. Subsequently, the classification vectors are processed by a multi-head classifier to generate probability distributions of $S$ codewords. The training objective is the multi-class Cross-Entropy, called the visual prediction loss:
@@
where $s_{m}^{*}$ denotes the ground-truth codeword index corresponding to the $m$-th classification head. During inference, the codewords with the highest probability are selected as the predicted (virtual) visual features $\hat{F}_{v}$.
To enable gradient backpropagation, we utilize a Gumbel-Softmax relaxation:
@@ 4.5 Multimodal fusion and alignment
where $\tau$ is the temperature coefficient. By minimizing $\mathcal{L}_{align}$, the similarity of matched pairs within the bias $B$ (Eq. 18) is maximized, while that of mismatched pairs is minimized, ensuring that the Biased Cross-Attention module can distinguish between valid visual features and noise. syntax constraint algorithm
@@ 4.6 Decoder
18:  else if $w\in\mathcal{O}_{unary}$ then  $\mathcal{S}.\text{Push}(1,\Phi^{\prime})$ // Unary operators require 1 nodes
@@ 4.7 Optimization
The overall loss function is a weighted sum of the skeleton prediction loss $\mathcal{L}_{sp}$ (Eq. 21), the quantization loss $\mathcal{L}_{q}$ (Eq. 14), the visual prediction loss $\mathcal{L}_{vp}$ (Eq. 15), the consistency loss $\mathcal{L}_{cons}$ (Eq. 23), and the alignment loss $\mathcal{L}_{align}$ (Eq. 20):
$\mathcal{L}_{total}=\mathcal{L}_{sp}+\lambda(\mathcal{L}_{q}+\mathcal{L}_{vp}+\mathcal{L}_{cons}+\mathcal{L}_{align}).$ (24)
@@ 5 Data preprocessing
In this section, we detail the data preprocessing pipeline used for generating the training dataset for ViSymRe. This process is divided into three stages: First, generating diverse equation skeletons based on a binary tree; Second, sampling the constant in the skeleton to generate parsed equations; Finally, sampling points to generate the input datasets.
@@ 6.1 Training details
20:   $\mathcal{L}_{total}=\mathcal{L}_{sp}+\lambda(\mathcal{L}_{q}+\mathcal{L}_{vp}+\mathcal{L}_{cons}+\mathcal{L}_{align})$ // Calculate total loss by Eq. 24
@@ 6.2 Testing details
ViSymRe faces the challenge of out-of-distribution robustness due to the sparsity of the sampled datasets during training. To ensure that ViSymRe can adapt to extremely large or small numerical scales that may appear in the test dataset (such as in the field of astronomy and nuclear physics), we propose a median-based Adaptive Magnitude Scaling (AMS) algorithm, aiming to scale the test datasets to a scale familiar to ViSymRe. Specifically, consider a $d$-D test dataset. For $i$-th feature $x_{i}$, we determine a magnitude $\rho_{x_{i}}$ based on the median of its absolute values:
@@ 6.3 Benchmarks
    Low-dimensional benchmarks: The low-dimensional benchmarks include "Nguyen" [61], "Keijzer"[16], "Korns" [18], "Livermore" [41], "Jin" [13] and "Neat". These benchmarks contain equations with up to three variables. We randomly sample 200 points from the sampling range reported in H to generate the test datasets.
@@ 6.6 Results on low-dimensional benchmarks
Table 3 compares the testing time of models. Observed that ViSymRe achieves the shortest testing time, outperforming the standard pre-trained baseline NeSymRes by a factor of four. Besides the convergence level, a possible explanation is that the syntax-constraint algorithm promotes ViSymRe to generate simpler equation skeletons, enabling faster constant optimization, thereby triggering early stopping more effectively than NeSymReS. Furthermore, RILS-ROLS requires less testing time than PYSR and uDSR, reflecting that the evolution of equations via perturbation may be more controllable when dealing with low-complexity scenarios than population-based methods.
@@
To evaluate the generalization ability of the equations generated by the models, we conducted extrapolation tests. Specifically, we construct the out-of-domain datasets by doubling the sampling range of the in-domain datasets. All models generate equations based on in-domain datasets and are tested for extrapolation on out-of-domain datasets. To demonstrate the superiority of the SR model in extrapolation scenarios, three classic machine learning models, including Multilayer Perceptron (MLP), Random Forest (RF), and Support Vector Machine (SVM), are introduced as supplementary baselines (detailed parameter settings are provided in LABEL:PYSR_UDSR_ML_parameters). To enhance experimental realism, we introduce noise at a level of 0.01 to the in-domain datasets. Fig. 6 presents the in-domain error distributions and out-of-domain error distributions of models. It can be observed that while the classic machine learning baselines achieve outstanding numerical approximation results on the in-domain datasets, their out-of-domain errors increased significantly. This phenomenon highlights the well-known tendency of them to overfit the local distribution of in-domain datasets rather than learning global laws. In contrast, SR models demonstrate superior extrapolation robustness. ViSymRe exhibits low median and peak errors across both in-domain and out-of-domain datasets, a significant achievement given the noisy condition. While other SR models also outperform classic machine learning models, they tend to minimize error by overfitting noise, thereby compromising their extrapolation potential, as evidenced by median in-domain errors that are significantly higher than median out-of-domain errors.
@@ 6.8 Results on SRSD-Feynman benchmark
To assess the ViSymRe’s performance in realistic scientific discovery scenarios, we evaluate it on the Phenomenological & First-principles problems. As illustrated in Fig. 11, ViSymRe demonstrates a superior trade-off between $R^{2}$ score and Complexity. ViSymRe achieves a median $R^{2}$ score close to $1.0$ with a remarkably tight distribution. Importantly, the advantage of ViSymRe lies in its interpretability, as reflected in the Complexity. While PYSR and RILS-ROLS achieve a high $R^{2}$ score, they exhibit high variance in Complexity, often yielding complex equations. Furthermore, ViSymRe produces fewer outliers in both $R^{2}$ score and Complexity compared to baselines, further corroborating its robustness when handling real-world noisy datasets. The details of the equations generated by ViSymRe are presented in Table 5
@@
Dataset equation $R^{2}$
bode $0.15e^{0.68x_{1}}+0.39$ 0.997
hubble $480x_{1}+169\sin{\left(10x_{1}+9.11\right)}+25$ 0.925
kepler $362x_{1}^{1.5}$ 0.999
ideal_gas $\log{\left(\frac{10.5x_{1}x_{2}}{x_{3}}-6.93\right)}$ 0.998
leavitt $0.43x_{1}+11.8+3.14e^{-0.32x_{1}^{3}}$ 0.980
rydberg $\frac{3.55x_{1}-3.73}{x_{2}^{0.69}}-15.9$ 0.993
absorption $0.0045x_{1}+3.57-3.43e^{-0.032x_{1}^{1.41}}$ 0.995
supernovae_zg $0.94-0.93e^{-\frac{89.4}{x_{1}^{2}}}$ 0.981

### v4

@@ Abstract
Extracting interpretable equations from observational datasets to describe complex natural phenomena is one of the core goals of artificial intelligence. This field is known as symbolic regression (SR). In recent years, Transformer-based paradigms have become a new trend in SR, addressing the well-known problem of inefficient search. However, the modal heterogeneity between datasets and equations often hinders the convergence and generalization of these models. In this paper, we propose ViSymRe, a Vision Symbolic Regression framework, to explore the positive role of visual modality in enhancing the performance of Transformer-based SR paradigms. To overcome the challenge where the visual SR model is untrainable in high-dimensional scenarios, we present Multi-View Random Slicing (MVRS). By projecting multivariate equations into 2-D space using random affine transformations, MVRS avoids common defects in high-dimensional visualization, such as variable degradation, non-linear interaction missing, and exponentially increasing sampling complexity, enabling ViSymRe to be trained with low computational costs. To support dataset-only deployment of ViSymRe, we design a dual-vision pipeline architecture based on generative techniques, which reconstructs visual features directly from the datasets via an auxiliary Visual Decoder and automatically suppresses the attention weights of reconstruction noise through a proposed Biased Cross-Attention feature fusion module, ensuring that subsequent processes are not affected by noisy modalities. Ablation studies demonstrate the positive contribution of visual modality to improving model convergence level and enhancing various SR metrics. Furthermore, evaluation results on mainstream benchmarks indicate that ViSymRe achieves competitive performance compared to baselines, particularly in low-complexity and rapid-inference scenarios.
@@ 1 Introduction
Recent advances in SR have introduced Transformer architectures [3, 64, 15, 66, 5], reformulating the SR problem as a dataset-to-equation translation task, similar to machine translation models. Different from traditional GP and DL-based models, Transformer-based paradigms leverage the attention mechanism to capture dependencies within datasets, enabling direct equation generation without iteration. However, despite improving deployment efficiency, these models often fail to adequately account for the modal gap between datasets and equations, which increases convergence difficulty. To address this issue, MMSR [31] pioneered multimodal learning in SR. It considers the fusion representation of the datasets and equations as direct inputs during training, and introduces multimodal techniques such as Cross-Attention feature fusion and alignment loss, demonstrating excellent performance compared to dataset-only baselines across several scenarios. Our solution further advances multimodal learning by incorporating visual modality, a readily available and key resource, but has been overlooked in previous research.
Currently, visual multimodal models have proven effective for enhancing various machine learning tasks, especially in machine translation. [11, 56]. The rationale is intuitive. For example, in machine translation, the English word “Dog” and the German word “Hund” share the same underlying visual perception. Analogously, in SR, although datasets and equations present different forms, they correspond to the same geometric topology.
@@
Based on MVRS, we propose ViSymRe, a Transformer-based Vision Symbolic Regression framework. While MVRS enables low-cost visualization, its application is limited by the analytical equations, a condition that holds only during training. To achieve dataset-only inference, ViSymRe employs an auxiliary Visual Decoder to predict visual features directly from dataset embeddings, thereby eliminating the need for real visual input. Diverging from traditional pixel-level reconstruction or direct feature regression, we incorporate a Vector Quantization mechanism to project visual features into a finite space [65]. Specifically, ViSymRe maintains a learnable Codebook that discretizes the continuous visual feature space into a finite set of codewords. The Visual Decoder performs multi-head classification conditioned on the dataset to predict the indices of codewords corresponding to the visual features. By transforming the visual reconstruction task from regression to classification, we avoid the common mean collapse issue in high-dimensional regression [38]. In essence, ViSymRe establishes a dual-visual pipeline: a real pipeline and a virtual pipeline. The former supervises the convergence of the Visual Encoder to ensure the validity of visual features, while the latter enables dataset-only inference.
@@
    We propose ViSymRe, a Transformer-based framework that integrates the visual modality as a bridge into SR to address the challenge of modal heterogeneity between datasets and equations faced by traditional pre-trained models. Extensive evaluations demonstrate that the proposed architecture is robust across various scenarios, particularly showing strong potential for low-complexity and rapid-inference tasks.
@@
    We propose Multi-View Random Slicing (MVRS), a visualization strategy to support large-scale pre-training. MVRS projects multivariate equations into 2-D space without causing variable degeneracy or exponential sampling complexity, thereby generalizing ViSymRe to high-dimensional scenarios at a low computational cost.
@@
    We construct a dual-visual pipeline that introduces virtual vision to reconcile conflicts between multimodal training and dataset-only inference. By employing a Biased Cross-Attention feature fusion module to suppress attention weights assigned to noisy components in the virtual vision, the two vision pipelines are ensured to converge synchronously.
@@ 2 Motivation
Transformer-based paradigms have become a new trend for SR. They enable one-time inference, thus avoiding expensive iterative search. However, adapting the machine translation paradigms to SR remains challenging. A key difficulty lies in the modal heterogeneity between the datasets and equations, which limits the models’ convergence. To address this problem, Multimodal learning applied to SR has been preliminarily explored in [31], ensuring modal homogeneity by treating the equations as input. Our solution is to introduce the visual modality as a bridge to alleviate this limitation. The dataset-only paradigm requires models to generate equations from local, fragmented data points, whereas the multimodal paradigm provides a global perspective.
@@ 3 Related work
Overall, Transformer-based models offer an advantage in inference efficiency, avoiding the iterative search of GP and DL-based models by leveraging priors learned during training. However, these models face challenges in out-of-distribution generalization. To bridge this gap, some rectification strategies, such as TPSR [57], SNR [34], and SR-GPT [30], have been proposed to offer feedback about fitting losses and complexity to generation policies. These rectification strategies are crucial for enhancing the generalization of Transformer-based models, significantly broadening their application scenarios at low cost.
@@ 4.2 Multi-view random slicing
Figure 2: Visualization results comparison of four visualization methods for the interaction term $x_{1}(x_{2}-x_{3})$. The proposed method (a) preserves the saddle shape via random isotropic slicing. Marginal projection (b) suffers from feature collapse, degenerating the non-linear interaction into a misleading linear response. Traditional methods (c and d): PCA exhibits information loss due to projection overlap, while t-SNE fails to reconstruct the manifold continuity, resulting in topological tearing even with dense sampling. Notably, MVRS requires only $N\cdot d$ points, while PCA and t-SNE require $N^{d}$ points. The computational burden is a key reason why traditional methods cannot support large-scale pre-training.
@@ 4.4 Visual Decoder
The Visual Decoder establishes a mapping between the dataset and Codebook, enabling dataset-only inference. This module operates essentially as a parallel classifier. We initialize $M$ learnable classification tokens $[CLS]s$, corresponding to the $M$ visual features output by the Visual Encoder. $[CLS]s$ aggregates with the dataset features via a Cross-Attention module to generate $M$ classification vectors. Subsequently, these classification vectors are processed by a multi-head classifier to generate probability distributions of $S$ codewords. The training objective is the multi-class Cross-Entropy, called the visual prediction loss:
@@
where $s_{m}^{*}$ denotes the ground-truth codeword index corresponding to the $m$-th classification head. The codewords with the highest probability are selected as the predicted (virtual) visual features $\hat{F}_{v}$.
To enable gradient backpropagation during training, we utilize a Gumbel-Softmax relaxation:
@@ 4.5 Multimodal fusion and alignment
where $\tau$ is the temperature coefficient. By minimizing $\mathcal{L}_{align}$, the similarity of matched pairs within the bias $B$ (Eq. 18) is maximized, while that of mismatched pairs is minimized, ensuring that the Biased Cross-Attention module can distinguish between valid visual features and noise.
@@ 4.6 Decoder
18:  else if $w\in\mathcal{O}_{unary}$ then  $\mathcal{S}.\text{Push}(1,\Phi^{\prime})$ // Unary operators require 1 node
@@ 4.7 Optimization
The overall loss function is a weighted sum of the quantization loss $\mathcal{L}_{q}$ (Eq. 14), the visual prediction loss $\mathcal{L}_{vp}$ (Eq. 15), the alignment loss $\mathcal{L}_{align}$ (Eq. 20), the skeleton prediction loss $\mathcal{L}_{sp}$ (Eq. 21), and the consistency loss $\mathcal{L}_{cons}$ (Eq. 23):
$\mathcal{L}_{total}=\mathcal{L}_{sp}+\lambda(\mathcal{L}_{q}+\mathcal{L}_{vp}+\mathcal{L}_{align}+\mathcal{L}_{cons}).$ (24)
@@ 5 Data preprocessing
In this section, we detail the data preprocessing pipeline used for generating the training dataset for ViSymRe. This process is divided into three stages: First, generating diverse equation skeletons based on a binary tree; Second, sampling the constant in the skeleton to generate analytical equations; Finally, sampling points to generate the input datasets.
@@ 6.1 Training details
20:   $\mathcal{L}_{total}=\mathcal{L}_{sp}+\lambda(\mathcal{L}_{q}+\mathcal{L}_{vp}+\mathcal{L}_{align}+\mathcal{L}_{cons})$ // Calculate total loss by Eq. 24
@@ 6.2 Testing details
ViSymRe faces the challenge of out-of-distribution robustness due to the sparsity of the training datasets. To ensure that ViSymRe can adapt to extremely large or small numerical scales that may appear in the test datasets (such as in the field of astronomy and nuclear physics), we propose a median-based Adaptive Magnitude Scaling (AMS) algorithm, aiming to scale the test datasets to a scale familiar to ViSymRe. Specifically, consider a $d$-D test dataset. For $i$-th feature $x_{i}$, we determine a magnitude $\rho_{x_{i}}$ based on the median of its absolute values:
@@ 6.3 Benchmarks
    Low-dimensional benchmarks: The low-dimensional benchmarks include "Nguyen" [63], "Keijzer"[17], "Korns" [19], "Livermore" [43], "Jin" [14] and "Neat". These benchmarks contain equations with up to three variables. We randomly sample 200 points from the sampling range reported in LABEL:tab:Details_on_low-dimensional_benchmarks to generate the test datasets.
@@ 6.6 Results on low-dimensional benchmarks
Table 3 compares the testing time of models. Observed that ViSymRe achieves the shortest testing time, outperforming the standard pre-trained baseline NeSymRes by a factor of four. Besides the convergence level, a possible explanation is that the syntax constraint algorithm promotes ViSymRe to generate simpler equation skeletons, enabling faster constant optimization, thereby triggering early stopping more effectively than NeSymReS. Furthermore, RILS-ROLS requires less testing time than PYSR and uDSR, reflecting that the evolution of equations via perturbation may be more controllable when dealing with low-complexity scenarios than population-based methods.
@@
To evaluate the generalization ability of the equations generated by the models, we conducted extrapolation tests. Specifically, we construct the out-of-domain datasets by doubling the sampling range of the in-domain datasets. All models generate equations based on in-domain datasets and are tested for extrapolation on out-of-domain datasets. To demonstrate the superiority of the SR model in extrapolation scenarios, three classic machine learning models, including Multilayer Perceptron (MLP), Random Forest (RF), and Support Vector Machine (SVM), are introduced as supplementary baselines (detailed parameter settings are provided in Table 8). To enhance experimental realism, we introduce noise at a level of 0.01 to the in-domain datasets. Fig. 6 presents the in-domain error distributions and out-of-domain error distributions of models. It can be observed that while the classic machine learning baselines achieve outstanding numerical approximation results on the in-domain datasets, their out-of-domain errors increased significantly. This phenomenon highlights the well-known tendency of them to overfit the local distribution of in-domain datasets rather than learning global laws. In contrast, SR models demonstrate superior extrapolation robustness. ViSymRe exhibits low median and peak errors across both in-domain and out-of-domain datasets, a significant achievement given the noisy condition. While other SR models also outperform classic machine learning models, they tend to minimize error by overfitting noise, thereby compromising their extrapolation potential, as evidenced by median out-of-domain errors that are significantly higher than that of in-domain errors.
@@ 6.8 Results on SRSD-Feynman benchmark
To assess the ViSymRe’s performance in realistic scientific discovery scenarios, we evaluate it on the Phenomenological & First-principles problems. As illustrated in Fig. 11, ViSymRe demonstrates a superior trade-off between $R^{2}$ score and Complexity. ViSymRe achieves a median $R^{2}$ score close to $1.0$ with a remarkably tight distribution. Importantly, the advantage of ViSymRe lies in its interpretability, as reflected in the Complexity. While PYSR and RILS-ROLS achieve a high $R^{2}$ score, they exhibit high variance in Complexity, often yielding complex equations. Furthermore, ViSymRe produces fewer outliers in both $R^{2}$ score and Complexity compared to baselines, further corroborating its robustness when handling real-world noisy datasets.